(2x-1)(x+1)=4
Solve quadtic formula

Answer by stanbon(75887)

(2x-1)(x+1)=4
2x^2 + 2x -x -1 = 4
-----
2x^2 +x -5 = 0
-----
x = [-1 +- sqrt(1-4*2*-5)]/4
------
x = [-1 +- sqrt(41)]/4
===========================
Cheers,
Stan H.
==============